OpenStudy (anonymous):

1. Which of the following will have a negative result? A – 8 – (-10) B 12 – 14 C 7- (-3) 1. Select the word that best completes the statement. A. Always B. Sometime C. Never 2. Find the difference -24-5=? A. -29 B. 29 C. -19 D. 19 3. Find the difference -14-(-16) =? A. 30 B. -30 C. 2 D. -2 4. Which of the following is equivalent to (-2) – 5? A. -2+ (-5) B. -2+5 C. 2 + (-5) D. 2- (-5)

OpenStudy (igreen):

For the first one, just simplify them all. A. -8 - (-10) = -8 + 10 B. 12 - 14 C. 7 - (-3) = 7 + 3

OpenStudy (anonymous):

I cant do that know I need the answers now I only have 20 minutes!

OpenStudy (anonymous):

Texas chic 101 can you help me I need help now!

OpenStudy (anonymous):

your mom017 help me 2 please!

OpenStudy (anonymous):

B. B. A. C. A. These are the answers in order

OpenStudy (anonymous):

okay so there the right one I will get 100% right?

OpenStudy (anonymous):

yes its really easy

OpenStudy (texaschic101):

whats the one about the statement...what statement ?